Solve for X. .125 .0312 · Solve for X. .38R Reviewed and Revised: 3/13/2013 RS. Page 4 of 5.
Answer #1

d-125 P 0-312 Sin 1 OM OP oP= OM Sin160 32 o 125/2 =0 2267 Sunlfo PorON 2 0 8517 PN X= PN - 0-312 0-8517-0.312 = 05397 %3D 0-